OpenStudy (anonymous):

find the slope and the equation of the tangent line for y=x-cos2x at (0,-1)

OpenStudy (watchmath):

\(y'=1+2\sin(2x)\) So at \((0,-1)\) the slope is 1. Then the line is \(y=x+b\). Plugin \(x=0,y=-1\) we conclude that \(b=-1\). So \(y=x-1\).

myininaya (myininaya):

y'=1-2(-sin(2x))=1+2sin(2x) y'(0)=1+2sin(0)=1+2(0)=1 so the slope at (0,-1) is 1 y=mx+b equation for tangent line y=x+b since m=1 so (x,y)=(0,-1) -1=0+b b=-1 so y=x-1

OpenStudy (anonymous):

you are a stud!